OpenStudy (anonymous):

see, now i'm just really mad......emailed the TA and asked why my percent error was 25%, he responded and said i switched two numbers. now i'm doing it again and i should be getting 3.2*10^-13 and i'm getting 3.2*10^-10....and i converted mm to m......wtf!!!!???

OpenStudy (anonymous):

C=epsilon naught (A/d). epsilon = 8.854*10^-12 A=252.4mm^2 or 0.252m^2 and d=6.9 mm or .0069 m.... C = 8.854*10^-12*(.2524/.0069) gives me 3.24*10^-10....should be *10^-13 i seriously don't get what the hell happened.

OpenStudy (anonymous):

@ganeshie8 ?????

ganeshie8 (ganeshie8):

\[\large \mathrm{ 252.4~ mm^2 ~\ne~ 0.252 ~m^2}\]

ganeshie8 (ganeshie8):

\(1 mm = 10^{-3} m \) so \(1 mm^2 =(10^{-3} m)^2 = 10^{-6}~m^2 \)

OpenStudy (anonymous):

oh. so it's .000252? grr...this is why i shouldn't do homework when i already have a headache from nearly being driven off the road and falling 30 feet to my deaht...

OpenStudy (anonymous):

thanks, ganeshie

ganeshie8 (ganeshie8):

:)